OpenStudy (anonymous):

f(x,y)=e^(sinx)+(x^5)y+ln(1+y^2) find fx and fy

OpenStudy (amistre64):

f(x,y)=e^(sinx)+(x^5)y+ln(1+y^2) fx = cos(x) e^(sin(x)) + 5x^4 y fy = x^5 +2y/(1+y^2) if i see it right

OpenStudy (anonymous):

\[f(x,y)=e^sinx +x^5y +\ln(1+y^2)\]

OpenStudy (anonymous):

that's what it looks like, is that the way you saw it?

OpenStudy (amistre64):

enclose exponents in {} to keep it together

OpenStudy (amistre64):

\begin{array}c f(x,y)=e^{sinx}+x^5\ y+ln(1+y^2)\\\\ fx = cos(x) e^{sin(x)} + 5x^4\ y \\\\ fy = x^5 +\frac{2y}{(1+y^2)} \\\\ \end{array}

OpenStudy (amistre64):

when doing partials; say like Fx, anything without an "x" is consider constant and is treated as such

OpenStudy (amistre64):

Fy same thing; treat everything that is not a y as a constant

OpenStudy (anonymous):

thank you so much :)

OpenStudy (amistre64):

youre welcome